# Weize Chen1  , Jiarui Yuan1∗, Chen Qian1Cheng Yang2 Zhiyuan Liu1, Maosong Sun1 1

### v1

@@ Abstract
Large Language Model (LLM) based multi-agent systems (MAS) show remarkable potential in collaborative problem-solving, yet they still face critical challenges: low communication efficiency, poor scalability, and a lack of effective parameter-updating optimization methods. We present Optima, a novel framework that addresses these issues by significantly enhancing both communication efficiency and task effectiveness in LLM-based MAS through LLM training. Optima employs an iterative generate, rank, select, and train paradigm with a reward function balancing task performance, token efficiency, and communication readability. We explore various RL algorithms, including Supervised Fine-Tuning, Direct Preference Optimization, and their hybrid approaches, providing insights into their effectiveness-efficiency trade-offs. We integrate Monte Carlo Tree Search-inspired techniques for DPO data generation, treating conversation turns as tree nodes to explore diverse interaction paths. Evaluated on common multi-agent tasks, including information-asymmetric question answering and complex reasoning, Optima shows consistent and substantial improvements over single-agent baselines and vanilla MAS based on Llama 3 8B, achieving up to 2.8x performance gain with less than 10% tokens on tasks requiring heavy information exchange. Moreover, Optima’s efficiency gains open new possibilities for leveraging inference-compute more effectively, leading to improved inference-time scaling laws. By addressing fundamental challenges in LLM-based MAS, Optima shows the potential towards scalable, efficient, and effective MAS [^1] https://chenweize1998.github.io/optima-project-page.
@@ 1 Introduction
Current LLM-based MAS face significant difficulties in meeting these challenges. The coordination and communication between agents often lack efficiency, resulting in verbose exchanges that lead to increased token usage, longer inference times, and higher computational costs (Li et al., 2024b). This inefficiency is exacerbated by the length bias inherent in LLMs due to alignment training (Saito et al., 2023; Dubois et al., 2024), which favors longer responses even when concise communication would suffice (Chen et al., 2024d). Moreover, while recent work has explored training LLMs for single-agent tasks (Song et al., 2024; Xiong et al., 2024) and MAS training is well-studied in reinforcement learning (Johnson et al., 2000; Lanctot et al., 2017; Baker et al., 2020), there remains a lack of parameter-updating methods specifically designed to optimize LLM-based MAS as a unified system. Existing approaches primarily rely on simple agent profile evolution (Chen et al., 2024b) or memory evolution (Qian et al., 2024a; b; Gao et al., 2024), which fail to address the core issues of communication efficiency and collective optimization.
@@
Can we develop a training framework that simultaneously enhances the communication efficiency and task effectiveness of LLM-based MAS? To address this question, we introduce Optima, an effective framework designed to optimize LLM-based MAS. At the heart of Optima is an iterative generate, rank, select, and train paradigm, incorporating a reward function that balances task performance, token efficiency, and communication interpretability. This approach enables the development of MAS that are not only effective and efficient but also maintain interpretable communication patterns. Based on the reward function, Optima leverages a combination of techniques to induce efficient and effective communication behaviors in LLM-based agents, including Supervised Fine-Tuning (SFT) (Zelikman et al., 2022; Gülçehre et al., 2023; Aksitov et al., 2023) and Direct Preference Optimization (DPO) (Rafailov et al., 2023; Pang et al., 2024), along with their hybrid variants. Furthermore, Optima introduces an integration of Monte Carlo Tree Search (MCTS)-inspired techniques for DPO data generation, conceptualizing conversation turns as tree nodes to explore diverse interaction trajectories efficiently.
@@
Importantly, by substantially reducing the number of tokens required for inference, Optima not only improves computational efficiency but also opens new possibilities for leveraging inference-compute more effectively. This reduction in token usage allows for more samples within the same computational constraints, potentially leading to better inference-time scaling laws. As recent work has shown the importance of inference-time compute in improving model performance (Wu et al., 2024; Brown et al., 2024; Chen et al., 2024a), Optima’s efficiency gains could be combined with techniques like majority voting (Wang et al., 2023), leading to more effective LLM systems.
@@
We evaluate Optima on a diverse set of tasks spanning two multi-agent settings: (a) information exchange, including information-asymmetric question answering (Chen et al., 2024d; Liu et al., 2024), and (b) debate, encompassing mathematical and reasoning tasks (Du et al., 2024; Chen et al., 2024b; Wu et al., 2023). Using Llama 3 8B (Meta, 2024) as our base model, we demonstrate that Optima consistently outperforms both single-agent MAS baselines, achieving up to 90% reduction in token usage and 2.8x increase in task performance.
@@
To summarize, our main contribution is Optima, a novel training framework that simultaneously optimizes communication efficiency and task effectiveness. To enhance high-quality training data generation in multi-agent settings for DPO, we introduce an integration of MCTS-like techniques. Our comprehensive empirical evaluation across diverse tasks demonstrates notable advancements in both token efficiency and task performance, while also providing insights into the learned communication patterns. Additionally, we examine the implications of Optima’s efficiency gains for inference-time scaling laws, underscoring its potential to improve the overall capabilities of LLM systems by enabling more effective utilization of inference-compute. By addressing the dual challenges of communication efficiency and collective optimization, our work underscores the importance of developing advanced training frameworks for LLM-based MAS and highlights efficiency as a crucial metric to consider. We believe Optima provides a solid foundation for future investigations into scaling and improving MAS and even general LLM systems.
@@ 2.1 Overview
$R(\tau_{i}^{j})=R_{\text{task}}(\tau_{i}^{j})-\lambda_{\text{token}}R_{\text{token}}(\tau_{i}^{j})+\lambda_{\text{loss}}\frac{1}{R_{\text{loss}}(\tau_{i}^{j})}.$ (1)
Here, $R_{\text{task}}:\mathcal{T}\rightarrow\mathbb{R}$ is the task-specific performance metric, $R_{\text{token}}(\tau_{i}^{j})=\frac{\#\text{Tokens}(\tau_{i}^{j})}{\max_{k}(\{\#\text{Tokens}(\tau_{i}^{k})\}_{k})}$ is the normalized token count, and $R_{\text{loss}}(\tau_{i}^{j})=g\big{(}\mathcal{L}(\mathcal{M}_{\text{base}},d_{i},\tau_{i}^{j})\big{)}$ is based on the language modeling loss of the base model $\mathcal{M}_{\text{base}}$, which we detail in Section E.2. The positive coefficients $\lambda_{\text{token}}$ and $\lambda_{\text{loss}}$ are hyper-parameters . This reward function is designed to balance multiple objectives simultaneously: $R_{\text{task}}$ ensures that the model improves on the intended task, $R_{\text{token}}$ encourages communication efficiency by penalizing verbose exchanges, and $R_{\text{loss}}$ regularizes language naturalness and readability by favoring trajectories that are probable under the base model. By incorporating these components, we aim to develop LLM-based MAS that are not only effective in their designated tasks but also efficient in their communication, while maintaining interpretability in their outputs, unlike the often incomprehensible communication in prior RL research (Lazaridou et al., 2017; Evtimova et al., 2018; Chaabouni et al., 2022).
@@
2.2 Initialization: Diversifying Agent Communication
Before starting the iterative training process, we address a critical challenge in LLM-based MAS: agents often produce responses in a similar style across conversation trajectories, even with high-temperature sampling. This homogeneity limits the exploration of diverse communication strategies, potentially hindering the optimization toward more efficient and effective interactions. Following the observation from AutoForm (Chen et al., 2024d), where LLMs can be explicitly prompted to leverage different more concise formats to communicate or reason without much compromise in performance, we introduce an initialization step that promotes diversity in agent communication.
Our approach leverages a pool of format specification prompts, $\mathcal{P}=\{p_{1},p_{2},...,p_{K}\}$, where each $p_{k}$ is a string specifying a particular response format (e.g., JSON, list, see Appendix F for concrete examples and creation process). For each task instance $d_{i}\in\mathcal{D}$, we generate $N$ conversation trajectories, each with a randomly selected format specification appended to the input task:
$\tau_{i}^{j}=\mathcal{M}_{\text{base}}(d_{i}\oplus p_{k_{j}}),\quad k_{j}\sim\text{Uniform}(1,K),\quad j=1,...,N,$ (2)
@@ 2.2 Initialization: Diversifying Agent Communication
We then evaluate these trajectories using the reward function defined in Eq. 1, for each $d_{i}$, we select the trajectory with the highest reward: $\tau_{i}^{*}=\operatorname*{arg\,max}_{j}R(\tau_{i}^{j})$. Finally, we select top k trajectories that exceed a predefined performance threshold $\theta_{\text{init}}$, resulting in a high-quality dataset:
$\mathcal{D}_{0}^{*}=\text{TopK}(\{(d_{i},\tau_{i}^{*})|R_{\text{task}}(\tau_{i}^{*})>\theta_{\text{init}},\forall d_{i}\in\mathcal{D}\},0.7|D|).$ (3)
Crucially, we remove the format specification prompts from the selected trajectories, resulting in a dataset of diverse, high-quality conversations without explicit format instructions. Using this dataset, we fine-tune the base model and obtain $\mathcal{M}_{\text{base}}$ to obtain $\mathcal{M}_{0}=\text{SFT}(\mathcal{M}_{\text{base}},\mathcal{D}_{0}^{*})$, which serves as the starting point for Optima, able to generate diverse communication patterns without explicit format prompting. We provide pseudo-code in Appendix B for better understanding. This initialization sets the stage for more effective exploration and optimization in the subsequent iterative training process.
Algorithm 1 Iterative Supervised Fine-Tuning
2:Optimized model $\mathcal{M}_{T}$
4:for $t=0$ to $T-1$ do
5:     $\mathcal{D}_{t}^{*}\leftarrow\emptyset$
6:     for each $d_{i}\in\mathcal{D}$ do
9:         if $R(\tau_{i}^{*})>\theta_{\text{sft}}$ then
10:              $\mathcal{D}_{t}^{*}\leftarrow\mathcal{D}_{t}^{*}\cup\{(d_{i},\tau_{i}^{*})\}$
11:         end if
12:     end for
15:end for
16:return $\mathcal{M}_{T}$
2.3 Framework Instantiation 1: Iterative Supervised Fine-Tuning
We introduce iterative Supervised Fine-Tuning (iSFT) as our first instantiation of Optima. At each iteration $t$, iSFT follows the same general procedure outlined in Algorithm 3, generating a set of $N$ conversation trajectories for each task training instance $d_{i}\in\mathcal{D}$ using the current model $\mathcal{M}_{t}^{\text{iSFT}}$. However, unlike initialization, iSFT omits the format specification pool, as $\mathcal{M}_{0}$ has already internalized diverse communication strategies. Unlike recent research on iterative training (Gülçehre et al., 2023; Aksitov et al., 2023), iSFT maintains a fixed reward threshold $\theta_{\text{SFT}}$ across iterations for data selection. After data generation, the model undergoes standard SFT. This process continues until a maximum number of iterations is reached. For clarity, the pseudo-code for iSFT is provided in Algorithm 1.
@@ 2.3 Framework Instantiation 1: Iterative Supervised Fine-Tuning
2.4 Framework Instantiation 2: Iterative Direct Preference Optimization
@@
Data Generation: To overcome these challenges, we integrate MCTS with DPO data collection for high-quality paired data generation in multi-agent settings. Our MCTS-based approach conceptualizes the multi-agent conversation as a tree, where nodes represent conversational turns, and edges represent continuations. This structure allows us to explore diverse interaction trajectories systematically and select high-quality paired data for DPO training. The MCTS process begins at the root node (initial task prompt) and proceeds as follows: (1) Expansion: We select a node to expand based on the following criteria. We first exclude leaf nodes and the second-to-last level nodes to avoid wasting computation on low-variance expansions, then exclude nodes with content similar to previously expanded nodes, measured based on edit distance (see Section E.1). From the remaining nodes, we select 10 nodes with the highest rewards and sample one using the softmax distribution over their rewards. (2) Simulation: For each selected node, we expand 3 trajectories, simulating the conversation to completion. (3) Backpropagation: Once a trajectory is completed and rewarded with Eq. 1, we update the estimated rewards of all nodes in the trajectory with the average rewards from their children. (4) Iteration: We repeat the above process 8 times, resulting in 24 trajectories. More iterations could potentially lead to more diverse and better-quality data.
@@
The iDPO process then proceeds iteratively, alternating between MCTS-based data generation and model updates using DPO. The pseudo-code for our iDPO process is presented in Algorithm 2.
Algorithm 2 Iterative Direct Preference Optimization
2:Optimized model $\mathcal{M}_{T}$
4:for $t=0$ to $T-1$ do
5:     $\mathcal{D}_{t}^{\text{DPO}}\leftarrow\emptyset$
6:     for each $d_{i}\in\mathcal{D}$ do
9:     end for
11:end for
12:return $\mathcal{M}_{T}$
2.5 Framework Instantiation 3: Hybrid Iterative Training
@@ 3 Experiments
Datasets. We evaluate Optima on two multi-agent settings: information exchange (IE) and debate. For IE, we use HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (2WMHQA) (Ho et al., 2020), TriviaQA (Joshi et al., 2017), and CBT (Hill et al., 2016). For multi-hop datasets (HotpotQA, 2WikiMultiHopQA), we split relevant contexts between two agents, ensuring the answer can only be deduced from information exchange. For TriviaQA and CBT, contexts are randomly assigned, challenging agents to identify and communicate the relevant information effectively. The debate setting employs GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), ARC’s challenge set (ARC-C) (Bhakthavatsalam et al., 2021) and MMLU (Hendrycks et al., 2021a), with one agent as solver and another as critic (Chen et al., 2024b). We use 0-shot for all benchmarks.
@@
Baselines. We compare against single-agent approaches: Chain-of-Thought (CoT) (Wei et al., 2022) and Self-Consistency (SC) with majority voting (Wang et al., 2023) on $n=8$ samples. Given that the generated responses for IE tasks are in free form, direct adaptation to majority voting is impractical. Therefore, we first compute the pairwise F1 score among the sampled answers, grouping those with a pairwise F1 score exceeding 0.9, and report the average F1 score against the label for all the answers in the largest grouping. In the multi-agent context, we compare against Multi-Agent Debate (MAD) from Du et al. (2024) and AutoForm (Chen et al., 2024d). MAD utilizes natural language for inter-agent communication, providing a baseline for common multi-agent dialogue, while AutoForm encourages agents to leverage concise, non-natural-language formats to achieve a better performance-cost ratio, offering a comparison point for efficiency-oriented MAS.
Training Setups. We use Llama 3 8B (Meta, 2024) as our base model across all benchmarks. Our experiments focus on two-agent scenarios without external tools, a design choice that allows us to isolate and analyze the core aspects of multi-agent communication and collaboration. By constraining our initial investigation to these fundamental settings, we can more clearly demonstrate the efficacy of Optima in optimizing inter-agent communication and task performance. This approach also provides a strong baseline for future research exploring more complex scenarios with multiple agents and tool use. Besides, we train a single model for both agents, although training separate models might yield improved performance, we leave it for future exploration. Detailed training configurations and prompts are provided in Appendices E and F.

### v2

# Weize Chen1, Jiarui Yuan1∗, Chen Qian1,Cheng Yang2, Zhiyuan Liu1, Maosong Sun1 1

@@ Abstract
Large Language Model (LLM) based multi-agent systems (MAS) show remarkable potential in collaborative problem-solving, yet they still face critical challenges: low communication efficiency, poor scalability, and a lack of effective parameter-updating optimization methods. We present Optima, a novel framework that addresses these issues by significantly enhancing both communication efficiency and task effectiveness in LLM-based MAS through training. Optima employs an iterative generate, rank, select, and train paradigm with a reward function balancing task performance, token efficiency, and communication readability. We explore various algorithms, including Supervised Fine-Tuning, Direct Preference Optimization, and their hybrid approaches, providing insights into their effectiveness-efficiency trade-offs. We integrate Monte Carlo Tree Search-inspired techniques for DPO data generation, treating conversation turns as tree nodes to explore diverse interaction paths. Evaluated on common multi-agent tasks, including information-asymmetric question answering and complex reasoning, Optima shows consistent and substantial improvements over single-agent baselines and vanilla MAS based on Llama 3 8B / 3.2 3B, achieving up to 2.8x performance gain with less than 10% tokens on tasks requiring heavy information exchange. Moreover, Optima’s efficiency gains enable more effective compute utilization during inference, leading to improved inference-time scaling laws. By addressing fundamental challenges in LLM-based MAS, Optima shows the potential towards scalable, efficient, and effective MAS.
Optima: Optimizing Effectiveness and Efficiency for LLM-Based Multi-Agent System
[…]
@@ 1 Introduction
Current LLM-based MAS face significant difficulties in meeting these challenges. The coordination and communication between agents often lack efficiency, resulting in verbose exchanges that lead to increased token usage, longer inference times, and higher computational costs (Li et al., 2024b). This inefficiency is exacerbated by the length bias inherent in LLMs due to alignment training (Saito et al., 2023; Dubois et al., 2024), which favors longer responses even when concise communication would suffice (Chen et al., 2024c). Moreover, while recent work has explored training LLMs for single-agent tasks (Song et al., 2024; Xiong et al., 2024) and MAS training is well-studied in reinforcement learning (Johnson et al., 2000; Lanctot et al., 2017; Baker et al., 2020), there remains a lack of parameter-updating methods specifically designed to optimize LLM-based MAS as a unified system. Existing approaches primarily rely on simple agent profile evolution (Chen et al., 2024b) or memory evolution (Qian et al., 2024a, b; Gao et al., 2024), which fail to address the core issues of communication efficiency and collective optimization.
@@
Can we develop a training framework that simultaneously enhances the communication efficiency and task effectiveness of LLM-based MAS? To address this question, we introduce Optima, an effective framework designed to optimize LLM-based MAS. At the heart of Optima is an iterative generate, rank, select, and train paradigm, incorporating a reward function that balances task performance, token efficiency, and communication readability. This approach enables the development of MAS that are not only effective and efficient but also maintain interpretable communication patterns. Based on the reward function, Optima leverages a combination of techniques to induce efficient and effective communication behaviors in LLM-based agents, including Supervised Fine-Tuning (SFT) (Zelikman et al., 2022; Gülçehre et al., 2023; Aksitov et al., 2023) and Direct Preference Optimization (DPO) (Rafailov et al., 2023; Pang et al., 2024), along with their hybrid variants. Furthermore, Optima introduces an integration of Monte Carlo Tree Search (MCTS)-inspired techniques for DPO data generation, conceptualizing conversation turns as tree nodes to explore diverse interaction trajectories efficiently.
@@
Importantly, by substantially reducing the number of tokens required for inference, Optima not only improves computational efficiency but also opens new possibilities for leveraging inference compute more effectively. This reduction in token usage allows for more samples within the same computational constraints, potentially leading to better inference-time scaling laws. As recent work has shown the importance of inference-time compute in improving model performance (Wu et al., 2024; Brown et al., 2024; Chen et al., 2024a), Optima’s efficiency gains could be combined with techniques like majority voting (Wang et al., 2023), leading to more effective LLM systems.
@@
We evaluate Optima on a diverse set of tasks spanning two multi-agent settings: (a) information exchange, including information-asymmetric question answering (Chen et al., 2024c; Liu et al., 2024), and (b) debate, encompassing mathematical and reasoning tasks (Du et al., 2024; Chen et al., 2024b; Wu et al., 2023). Using Llama 3 8B / 3.2 3B (Meta, 2024) as our base model, we demonstrate that Optima consistently outperforms both single-agent MAS baselines, achieving up to 90% reduction in token usage and 2.8x increase in task performance.
@@
To summarize, our main contribution is Optima, a novel training framework that simultaneously optimizes communication efficiency and task effectiveness. To enhance high-quality training data generation in multi-agent settings for DPO, we introduce an integration of MCTS-like techniques. Our comprehensive empirical evaluation across diverse tasks demonstrates notable advancements in both token efficiency and task performance, while also providing insights into the learned communication patterns. Additionally, we examine the implications of Optima’s efficiency gains for inference-time scaling, underscoring its potential to improve the LLM systems by enabling more effective utilization of inference-compute. By addressing the dual challenges of communication efficiency and collective optimization, our work underscores the importance of developing advanced training frameworks for LLM-based MAS and highlights efficiency as a crucial metric to consider. We believe Optima provides a solid foundation for future investigations into scaling and improving MAS and general LLM systems.
@@ 2.1 Overview
$R(\tau_{i}^{j})=R_{\text{task}}(\tau_{i}^{j})-\lambda_{\text{token}}R_{\text{token}}(\tau_{i}^{j})+\lambda_{\text{loss}}\frac{1}{R_{\text{loss}}(\tau_{i}^{j})}$.
Here, $R_{\text{task}}:\mathcal{T}\rightarrow\mathbb{R}$ is the task-specific performance metric, $R_{\text{token}}(\tau_{i}^{j})=\frac{\#\text{Tokens}(\tau_{i}^{j})}{\max_{k}(\{\#\text{Tokens}(\tau_{i}^{k})\}_{k})}$ is the normalized token count, and $R_{\text{loss}}(\tau_{i}^{j})=g\big{(}\mathcal{L}(\mathcal{M}_{\text{base}},d_{i},\tau_{i}^{j})\big{)}$ is based on the language modeling loss of the base model $\mathcal{M}_{\text{base}}$, which we detail in Section H.2. The positive coefficients $\lambda_{\text{token}}$ and $\lambda_{\text{loss}}$ are hyper-parameters . This reward function is designed to balance multiple objectives simultaneously: $R_{\text{task}}$ ensures that the model improves on the intended task, $R_{\text{token}}$ encourages communication efficiency by penalizing verbose exchanges, and $R_{\text{loss}}$ regularizes language naturalness and readability by favoring trajectories that are probable under the base model. By incorporating these components, we aim to develop LLM-based MAS that are not only effective in their designated tasks but also efficient in their communication, while maintaining interpretability in their outputs, unlike the often incomprehensible communication in prior RL research (Lazaridou et al., 2017; Evtimova et al., 2018; Chaabouni et al., 2022).
@@
2.2 Initialization
Before starting the iterative training process, we address a critical challenge in LLM-based MAS: agents often produce responses in a similar style across conversation trajectories, even with high-temperature sampling. This homogeneity limits the exploration of diverse communication strategies, potentially hindering the optimization toward more efficient and effective interactions. Following the observation from AutoForm (Chen et al., 2024c), where LLMs can be explicitly prompted to leverage different more concise formats to communicate or reason without much compromise in performance, we introduce an initialization step that promotes diversity in agent communication.
Our approach leverages a pool of format specification prompts, $\mathcal{P}=\{p_{1},p_{2},...,p_{K}\}$, where each $p_{k}$ is a string specifying a particular response format (e.g., JSON, list, see Appendix I for concrete examples and creation process). For each task instance $d_{i}\in\mathcal{D}$, we generate $N$ conversation trajectories, each with a randomly selected format specification appended to the input task:
$\tau_{i}^{j}=\mathcal{M}_{\text{base}}(d_{i}\oplus p_{k_{j}}),\ k_{j}\sim\text{Uniform}(1,K),$ (2)
@@ 2.2 Initialization
We then evaluate these trajectories using the reward function defined in Eq. 1, for each $d_{i}$, we select the trajectory with the highest reward: $\tau_{i}^{*}=\operatorname*{arg\,max}_{j}R(\tau_{i}^{j})$. Finally, we select top 70% trajectories that exceed a predefined performance threshold $\theta_{\text{init}}$, resulting in a high-quality dataset:
$\mathcal{D}_{0}^{*}=\text{TopK}(\big{\{}(d_{i},\tau_{i}^{*})\,\big{|}\,R_{\text{task}}(\tau_{i}^{*})>\theta_{\text{init}},\forall d_{i}\in\mathcal{D}\big{\}},70\%).$
Crucially, we remove the format specification prompts from the selected trajectories, resulting in a dataset of diverse, high-quality conversations without explicit format instructions. We then fine-tune the base model $\mathcal{M}_{\text{base}}$ to obtain $\mathcal{M}_{0}=\text{SFT}(\mathcal{M}_{\text{base}},\mathcal{D}_{0}^{*})$, which serves as the starting point for Optima, able to generate diverse communication patterns without explicit format prompting. We provide pseudo-code in Appendix B for better understanding. This initialization sets the stage for more effective exploration and optimization in the subsequent iterative training process.
2.3 Instantiation 1: Iterative SFT
We introduce iterative Supervised Fine-Tuning (iSFT) as our first instantiation of Optima. At each iteration $t$, iSFT follows the same general procedure outlined in Algorithm 1, generating a set of $N$ conversation trajectories for each task training instance $d_{i}\in\mathcal{D}$ using the current model $\mathcal{M}_{t}^{\text{iSFT}}$. However, unlike initialization, iSFT omits the format specification pool, as $\mathcal{M}_{0}$ has already internalized diverse communication strategies. Unlike recent research on iterative training (Gülçehre et al., 2023; Aksitov et al., 2023), iSFT maintains a fixed reward threshold $\theta_{\text{SFT}}$ across iterations for data selection. The model is then trained with standard SFT. This process continues until a maximum number of iterations is reached. For clarity, the pseudo-code for iSFT is provided in Appendix B.
@@ 2.3 Instantiation 1: Iterative SFT
2.4 Instantiation 2: Iterative DPO
@@
Data Generation: To overcome these challenges, we integrate MCTS with DPO data collection for high-quality paired data generation in multi-agent settings. Our MCTS-based approach conceptualizes the multi-agent conversation as a tree, where nodes represent conversational turns, and edges represent continuations. This structure allows us to explore diverse interaction trajectories systematically and select high-quality paired data for DPO training. The MCTS process begins at the root node (initial task prompt) and proceeds as follows: (1) Expansion: We select a node to expand based on the following criteria. We first exclude leaf nodes and the second-to-last level nodes to avoid wasting computation on low-variance expansions, then exclude nodes with content similar to previously expanded nodes, measured based on edit distance (see Section H.1). From the remaining nodes, we select 10 nodes with the highest rewards and sample one using the softmax distribution over their rewards. (2) Simulation: For each selected node, we expand 3 trajectories, simulating the conversation to completion. (3) Backpropagation: Once a trajectory is completed and rewarded with Eq. 1, we update the estimated rewards of all nodes in the trajectory with the average rewards from their children. (4) Iteration: We repeat the above process 8 times, resulting in 24 trajectories. More iterations could potentially lead to more diverse and better-quality data.
@@
The iDPO process then proceeds iteratively, alternating between MCTS-based data generation and model updates using DPO. The pseudo-code for our iDPO process is presented in Appendix B.
2.5 Instantiation 3: Hybrid Iterative Training
@@ 3 Experiments
Datasets. We evaluate Optima in two settings: information exchange (IE) and debate. For IE, we use HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (2WMHQA) (Ho et al., 2020), TriviaQA (Joshi et al., 2017), and CBT (Hill et al., 2016). For multi-hop datasets (HotpotQA, 2WMHQA), we split relevant contexts between two agents, ensuring the answer can only be deduced from information exchange. For TriviaQA and CBT, contexts are randomly assigned, challenging agents to communicate and identify the relevant information. The debate setting employs GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), ARC’s challenge set (ARC-C) (Bhakthavatsalam et al., 2021) and MMLU (Hendrycks et al., 2021a), with one agent as solver and another as critic (Chen et al., 2024b). We use 0-shot for all benchmarks.
@@
Baselines. We compare against single-agent approaches: Chain-of-Thought (CoT) (Wei et al., 2022) and Self-Consistency (SC) with majority voting (Wang et al., 2023) on $n=8$ samples. For IE tasks, direct majority voting is impractical due to free-form responses. Instead, we compute pairwise F1 scores, group answers with scores above 0.9, and report the average F1 score of the largest group against the label. In multi-agent settings, we compare against Multi-Agent Debate (MAD) (Du et al., 2024) and AutoForm (Chen et al., 2024c). MAD uses natural language for inter-agent communication, while AutoForm employs concise, non-natural-language formats for better performance-cost efficiency.
Training Setups. We use Llama 3 8B / 3.2 3B (Meta, 2024) as our base model, focusing on two-agent scenarios without external tools to isolate core multi-agent communication and collaboration. A single model is trained for both agents, with separate model training left for future work. Iterative training completes within 12 hours on 8 A100 GPUs for most tasks, except MATH, which takes around 24 hours. More details are in Appendices H and I.